OpenStudy (anonymous):

IF ANYONE COULD ANWSER THIS IT WOULD BE GREAT:D The data represents the miles per gallon (mpg) for vehicles manufactured by a company in the past five years. Find the inter-quartile range of the data. Data: 28.7, 25.5, 21.8, 21.6, 20.5, 16.7, 16.3, 15.4, 13.9. 6.4 7.8 5.5 10.1

OpenStudy (kropot72):

The first step is to order the data from lowest to highest values from left to right. Can you do that?

OpenStudy (anonymous):

yes! i can do that!

OpenStudy (anonymous):

13.9,15.4,16.3,16.7,20.5,21.6,21.8,25.5,28.7

OpenStudy (kropot72):

The median value (the middle data item) is 20.5. There are now two subsets as follows: {13.9, 15.4, 16.3, 16.7} and {21.6, 21.8, 25.5, 28.7}. The quartiles are the medians of these two subsets, which in these two cases is the value half way between the two either side of the middle. lower quartile = (15.4 + 16.3)/2= 15.85 upper quartile = (21.8 + 25.5)/2 = 23.65 The interquartile range is the difference between the upper and lower quartiles. Can you find it now?

OpenStudy (anonymous):

is it 7.8?

OpenStudy (kropot72):

You are correct!

OpenStudy (anonymous):

thanks sooo much!!

OpenStudy (kropot72):

You're welcome :)